# Lorentzian Graph Isomorphic Network

### v2

@@ Abstract
We introduce the Lorentzian Graph Isomorphic Network (LGIN), a novel graph neural network (GNN) designed to operate in hyperbolic spaces, leveraging the Lorentzian model to enhance graph representation learning. Existing GNNs primarily operate in Euclidean spaces, which can limit their ability to capture hierarchical and multi-relational structures inherent to complex graphs. LGIN addresses this by incorporating curvature-aware aggregation functions that preserve the Lorentzian metric tensor, ensuring embeddings remain constrained within the hyperbolic space by proposing a new update rule that effectively captures both local neighborhood interactions and global structural properties, enabling LGIN to distinguish non-isomorphic graphs with expressiveness atleast approximately as powerful as the Weisfeiler-Lehman test. Through extensive evaluations across nine benchmark datasets, including molecular and protein structures, LGIN consistently outperforms or matches state-of-the-art GNNs, demonstrating its robustness and efficacy in modeling complex graph structures. To the best of our knowledge, this is the first study to extend the concept of a powerful graph neural network to Riemannian manifolds, paving the way for future advancements in hyperbolic graph learning.
@@ 1 Introduction
However, extending such frameworks to the hyperbolic analogy is inherently challenging due to the need to ensure that embeddings remain constrained within the Riemannian manifold. In particular, ensuring that the operations performed during aggregation and transformation preserve the hyperbolic metric is non-trivial. Motivated by the potential to unify high expressive power with geometry suited for hierarchical data, we propose the Lorentzian Graph Isomorphic Network (LGIN). LGIN operates on the hyperboloid (Lorentzian) model of hyperbolic space. Our design prioritizes maintaining the intrinsic geometric properties and preserving the Lorentzian metric during aggregation and transformation, which presents challenges for achieving strictly injective message-passing. To address this, LGIN introduces a curvature-aware aggregation mechanism and update rule designed to approximate the expressive power of Euclidean GINs by carefully handling multisets within the hyperbolic geometry. We show that LGIN enhances discriminative capability compared to Euclidean GINs, particularly for graphs with underlying hierarchical or complex relational patterns, by integrating structural isomorphism testing principles with the geometric inductive biases of hyperbolic space by defining a theoretical framework for a powerfull GNN on hyperboli spaces.
@@
  • To the best of our knowledge, this is the first study that introduces an isomorphic Graph Neural Network(GNN) that incorporates both constant and variable curvature information.
  • We show that GNNs built using the hyperbolic-tangent-hyperbolic approach while preserving the metric tensor at the origin high discriminative capability.
  • We propose the theoretical framework for Lorentzian Graph Isomorphic Netork(LGIN), with a modified cardinality-scaled aggregation rule to preserve the cardinality information of multisets, as an approximately powerful alternative to Graph Isomorphic Network(GIN) and other state of the art hyperbolic graoph neural networks such as LGCN.
@@ 4 Proposed Framework
In this section, we define a powerful Graph Neural Network(GNN) in the hyperbolic space. We show that it can distinguish different graph structures by mapping them to different embeddings in the hyperboloid model and is as powerful as the WL test.
@@ 4.1 Overview
Theorem 4.1.
@@
We present the proof for theorem 4.1 in appendix A. Therefore, we can demonstrate that a graph neural network mapping node features onto a hyperboloid can achieve representational power approximately comparable to the Weisfeiler-Lehman (WL) test. Further, the following must be taken into consideration:
  • A powerful graph neural networks requires a strictly injective aggregation and readout functions(Xu et al., 2018). However, directly translating strictly injective operations to Riemannian manifolds while preserving geometric properties is non-trivial. Our goal is to design a network that achieves high expressive power by carefully approximating these properties within the Lorentzian geometry.
@@ 4.3 Lorentzian Graph Isomorphic Network
The pseudo code for the above rules is given in algorithm 1. Each variable and function is described in appendix C.
Algorithm 1 LorentzGIN Forward Pass
1:$manifold$, $eps$, $c\_in$, $nn$, $x$, $adj$
2:$x\leftarrow manifold.exp\_map\_zero(x,c\_in)$
3:if $x$ is Tensor then
4:     $x\leftarrow(x,x)$
5:end if
6:$out\leftarrow\text{aggregate}(x[0],adj)$
7:$x_{r}\leftarrow x[1]$
8:if $x_{r}\neq\text{None}$ then
9:     $log\_out\leftarrow manifold.log\_map\_zero(out,c=c\_in)$
10:     $log\_x\_r\leftarrow manifold.log\_map\_zero(x_{r},c=c\_in)$
11:     $pt\_xr\leftarrow manifold.ptransp(x_{r},out,log\_x\_r,c=c\_in)$
12:     $out\leftarrow manifold.exp\_map\_zero(log\_out+(1+eps)\times pt\_xr,c=c\_in)$
13:end if
14:return $nn(out)$
@@ 4.3.1 Justification of Lorentz Centroid
Equation 19 ensures embeddings reside on the hyperbolic manifold after aggregation i.e the preservation of metric tensors. Normalization prevents the node embeddings from collapsing into the origin or drifting too far. However, since attention-based aggregators are usually not strictly injective, it may lead to non-isomorphic graphs having the same embeddings. To mitigate this, we modify the Lorentz aggregation function to preserve the cardinality information of multisets during aggregation which can approximate the injective property(Zhang & Xie, 2020).
@@ 5.2.3 Performance on MoleculeNet Datasets
Table 3: Analysis of our framework’s performance on MoleculeNet baselines. The best score is marked in bold and the second best score is underlined. GIN models have been marked with an asterisk to indicate that our proposed model significantly outperforms GIN baselines. Hyphen indicates OOM on an 8GB M1 Macbook Pro.
@@ 5.2.4 Empirical Analysis on Curvature and Parallel Transport
Appendix C Variable Description of Algorithm 1
Algorithm 1 describes the update rule of LGIN. The following is the description of each variable in algorithm 1:
  • $manifold$ : Lorentzian manifold class containing all operations
  • $eps$ : Epsilon value controlling the importance of current node features
  • $c\_in$ : Input curvature
  • $nn$ : Lorentz Transformation Function
  • $x$ : Node Features
  • $adj$ : Adjacency Matrix
  • $out$ : Output of Lorentz Aggregation
  • $x_{r}$ : Node Features before aggregation
  • $log\_out$ : Logarithmic map of updated features after aggregation
  • $log\_x\_r$ : Logarithmic map of initial node features
  • $pt\_xr$ : Node Features adjusted after Parallel Transport
The functions for logarithmic, exponential maps and parallel transport have been described in section 3.3.
Appendix D Additional Related Work

### v3

# {textblock*}

@@ Abstract
While graph neural networks (GNNs) operating in hyperbolic spaces have shown promise for modeling hierarchical and complex relational data, a critical limitation often overlooked is their potentially limited discriminative power compared to their Euclidean counterparts or fundamental graph isomorphism tests like the Weisfeiler-Lehman (WL) hierarchy. Ex- isting hyperbolic aggregation schemes, while curvature-aware, may not sufficiently capture the intricate structural differences required to robustly distinguish non-isomorphic graphs owing to non-injective aggregation functions. To address this expressiveness gap in hyper- bolic graph learning, we introduce the Lorentzian Graph Isomorphic Network (LGIN), a novel GNN designed to achieve enhanced discriminative capabilities within the Lorentzian model of hyperbolic space. LGIN proposes a new update rule that effectively combines local neighborhood information with a richer representation of graph structure designed to preserve the Lorentzian metric tensor. This represents a significant step towards building more expressive GNNs in non-Euclidean geometries, overcoming a common bottleneck in current hyperbolic methods. We conduct extensive evaluations across nine diverse benchmark datasets, in- cluding molecular and protein structures. LGIN consistently outperforms or matches state- of-the-art hyperbolic and Euclidean GNNs, showcasing its practical efficacy and validating its superior ability to capture complex graph structures and distinguish between different graphs. To the best of our knowledge, LGIN is the first work to study the framework behind a powerful GNN on the hyperbolic space. The code for our paper can be found at https://github.com/Deceptrax123/LGIN
@@ 1 Introduction
However, extending such frameworks to the hyperbolic space is inherently challenging due to the need to ensure that embeddings remain constrained within the Riemannian manifold. In particular, ensuring that the operations performed during aggregation and transformation preserve the hyperbolic metric is non-trivial. Motivated by the potential to unify high expressive power with geometry suited for hierarchical data, we propose the Lorentzian Graph Isomorphic Network (LGIN). Our design prioritizes maintaining the intrinsic geometric properties and preserving the Lorentzian metric during aggregation and transformation, which presents challenges for achieving strictly injective message-passing. To effectively bridge this expressiveness gap, LGIN introduces a novel node update rule incorporating a cardinality-aware aggregation mechanism. This design is specifically crafted to approximate the strong discriminative capabilities associated with Euclidean Graph Isomorphism Networks (GINs) by preserving crucial information often lost in standard hyperbolic aggregations. Our empirical results demonstrate that LGIN achieves superior discriminative capabilities compared to existing hyperbolic GNNs, including LGCN and other architectures. We further provide a key contribution by empirically demonstrating that the proposed Hyperbolic-Tangent-Hyperbolic approach, when coupled with cardinality-aware aggregation, yields a GNN with expressivity at least approximately equivalent to the 1-dimensional Weisfeiler-Lehman test. By rigorously integrating structural isomorphism testing principles with the inherent geometric inductive biases of hyperbolic space, this work establishes a foundational theoretical framework for designing powerful and highly-discriminative GNNs on non-Euclidean manifolds.
@@
  • To the best of our knowledge, this is the first study that makes an attempt to study an isomorphic Graph Neural Network(GNN) that incorporates both constant and variable curvature information.
  • We show that GNNs built using the hyperbolic-tangent-hyperbolic approach while preserving the metric tensor at the origin has high discriminative capability.
  • We propose the theoretical framework for Lorentzian Graph Isomorphic Netork(LGIN), with a modified cardinality-scaled aggregation rule to preserve the cardinality information of multisets and a new update rule, as an approximately powerful alternative to Graph Isomorphic Network(GIN) and other state of the art hyperbolic graoph neural networks such as LGCN.
@@ 4 Proposed Framework
In this section, we discuss the framework for an approximately powerful GNN on the Lorentzian model. Note that this can be extended to other hyperbolic models as well through their respective exponential and logarithmic maps. We discuss the properties of the manifold that allow for discriminative GNNs, however such frameworks are limited by the need of preserving the metric tensor and attention aggregated weights. Throughout our study, we consider more importance to metric tensor preservation than injectivity of aggregation.
@@ 4.1 Overview
Theorem 4.1 (Manifold Property for powerful GNNs).
@@
We present the proof for theorem 4.1 in appendix A. Therefore, we can demonstrate that a graph neural network mapping node features onto a hyperboloid can achieve representational power approximately comparable to the Weisfeiler-Lehman (WL) test owing to the property of the metric embeddings on a hyperbolic manifold, thus describing a desirable embedding. However, the following must be taken into consideration:
  • A powerful graph neural networks requires strictly injective aggregation and readout functions(Xu et al., 2018). However, directly translating strictly injective operations to Riemannian manifolds while preserving geometric properties is non-trivial. Our goal is to design a network that achieves high expressive power by carefully approximating these properties within the Lorentzian geometry.
@@ 4.3.1 Justification of Lorentz Centroid
Equation 19 ensures embeddings reside on the hyperbolic manifold after aggregation i.e the preservation of metric tensors. Normalization prevents the node embeddings from collapsing into the origin or drifting too far. However, since attention-based aggregators are ot strictly injective, it may lead to non-isomorphic graphs having the same embeddings. To mitigate this, we modify the Lorentz aggregation function to preserve the cardinality information of multisets during aggregation which can approximate the injective property(Zhang & Xie, 2020).
@@ 5.2.3 Performance on MoleculeNet Datasets
Table 3: Analysis of our framework’s performance on MoleculeNet baselines. The best score is marked in bold and the second best score is underlined. GIN models have been marked with an asterisk to indicate that our proposed model significantly outperforms GIN baselines. Hyphen indicates OOM on an 8GB M2 Macbook Pro.
@@ 5.2.4 Empirical Analysis on Curvature and Parallel Transport
6 Conclusion and Limitations
In this paper, we address the need for simple yet effective graph neural networks on hyperbolic manifolds capable of capturing both semantic and structural features. We propose the Lorentzian Graph Isomorphic Network (LGIN), which, to the best of our knowledge, is the first framework to extend graph isomorphism tests to hyperbolic spaces with variable and fixed curvature. Through extensive experiments, we demonstrate that LGIN consistently matches or surpasses state-of-the-art models across multiple graph benchmarks. Additionally, we provide a detailed explanation of our update rule, including rigorous derivations for each step, thereby establishing a framework for an approximate solution for the expressivity bottleneck on hyperbolic spaces.
However, our analysis does not offer a comprehensive framework for hyperbolic graph neural networks that directly map between hyperbolic spaces without relying on tangent space projections. Further investigation is required to develop such transformations and fully explore the potential of these models, which we leave as a direction for future research.
@@ 6 Conclusion and Limitations
Appendix C Additional Related Work